# Friederike Becker1, Fabian Krüger1 and Melanie Schienle1,2

### v3

@@ 3.2 Choice of Tuning Parameters
Table 5 and Figure 8 in the supplement summarize the training sample performance for the different combinations of the tuning parameters. Based on the performance in the training sample, and the considerations in Section 2, we use a rolling window of $R=11$ observations, the absolute error method, and the default empirical quantile method (type = 7) [^1] Note that for $R=11$ and the absolute error method, this is also equivalent to the inverse of the empirical distribution function, that is, type = 1. As we do not estimate quantiles at the tails of the distribution the arguments of Hyndman and Fan (1996) in favor of the type = 8 method do not apply for our setting. of the quantile function in R. Our training sample results indicate that shorter window lengths $R<11$ perform worse in terms of the IS, see Figure 8 in the appendix. The difference in performance is particularly pronounced for very short window lengths $74\leq R<$ 7 and at longer forecast horizons. By contrast, the benefit of increasing $R$ from $8$ to $11$ is remarkably small in all empirical setups (absolute and directional errors, all forecast horizons). Hence the marginal benefit of increasing the sample size seems to vanish rapidly as $R$ increases. We exclude choices $R>11$ since we cannot evaluate their performance given the time span of our training data, while an extension of the training data would have reduced the time span of our holdout sample, thus limiting our ability to assess out-of-sample performance. From a statistical perspective, using a short window length $R$ ensures adaptability to potential structural breaks (c.f. Inoue et al., 2017). We find that coverage rates on the training data are generally more favorable for the absolute than for the directional error method, sometimes on the order of about 10 percentage points (see Table 5 in the supplement), while interval scores are similar. We additionally present illustrative insights from the training period in Section B.3.1 in the supplement. Albeit attractive in principle, we argue that directional forecast errors sometimes display undesirable practical properties in our setup, and are thus not a good candidate to account for a potential skew in the distribution of forecast errors.
@@ 3.3 Benchmark Models
In order to put the performance of our proposed method in perspective, we compare it to several benchmarks. For ease of presentation, we focus on three main benchmark models in our main empirical analysis that each represent one of the following broader model types: a) univariate or b) multivariate time series models yielding different point predictions, where prediction intervals are calculated either based on past performance forecast errors as in our approach or based on c) the respective time series model specification. For further variants of these benchmark models and respective comparisons we refer to Section 3.5.3 below as well as Section B of the supplement. In particular, Tables 6 and 7 of the supplement provide comprehensive results in terms of evaluation metrics.
All considered benchmarks are of (vector) autoregressive time series form, which is one of the most common choices in macroeconomic forecasting. For example, Faust and Wright (2009) show that simple univariate autoregressions provide good forecasts of U.S. GDP growth when equipped with an appropriate information set, and Carriero et al. (2024) argue that vector autoregressive models with stochastic volatility compare favorably to nonlinear quantile regression methods for U.S. GDP growth, inflation and unemployment. We fit the main benchmark models to quarterly data from the OECD for inflation and GDP growth (see Section B.1 in the supplement), using expanding windows of training data. In order to roughly match the timing of IMF forecasts, we equip the model-based spring forecasts with data up until the first quarter, and the model-based fall forecasts with data up until the third quarter. As we argue in Section 3.5.1, evaluation with respect to IMF data is still justified.
@@
For the second main benchmark (denoted BVAR), we consider a bivariate Bayesian vector autoregressive specification for inflation and GDP growth. As for the IMF forecasts (and for the AR model), we construct prediction intervals from empirical quantiles of the BVAR’s past forecast errors. Our BVAR implementation is based on the R package bvarsv (Krüger, 2015). While the latter considers the flexible model by Primiceri (2005) and Del Negro and Primiceri (2015), we restrict the priors in our main benchmark to shut off time variation in the model’s mean and covariance parameters, as in Knüppel et al. (2023, Section 5). We find that this simpler specification yields relatively stable parameter estimates for all countries and time periods. By contrast, Primiceri’s original priors that allow for time-varying parameters and stochastic volatility run into problems for some countries such as France and Great Britain after the extreme GDP growth rates observed at the onset of the Covid-19 pandemic. [^2] This problem could possibly be resolved by adjusting the model’s prior parameters to allow for an intermediate degree of flexibility, or by using a stochastic volatility specification that explicitly addresses Covid-19 outliers (as in Carriero et al., 2024). However, the details of such an adjustment would create many additional degrees of freedom in our setup, given that the severity of the Covid-19 outliers differs across the G7 countries. We hence use a simpler specification with constant parameters for our main analysis. Within the additional analyses discussed in Section 3.5.3, we also consider an approach that uses a stochastic volatility specification until 2019 and a constant volatility specification afterwards. This mixed variant entails a manual choice of regime switch. Even then, however, the performance of the model with constant parameters is largely similar with the exception of inflation at the next-year horizons where the mixed BVAR can outperform all methods and in particular our simple IMF-based approach.
@@ 3.4 Forecast Performance on the Hold-Out Data
We also consider Diebold and Mariano (1995) for the null hypothesis of equal predictive accuracy of the IMF-based method and each benchmark method. In order to limit the number of tests conducted, we average the interval scores across countries at each date (for a given variable, forecast horizon and benchmark model). This yields 24 tests in total (two variables, four horizons, and the AR, BVAR and BVAR-direct benchmark models). We consider two-sided tests, a significance level of 95% and use the R package sandwich package and its function NeweyWest (Zeileis, 2004; Zeileis et al., 2020) for the autocorrelation-robust variance estimator that enters the test statistic. Among the 24 test setups, the null of equal predictive accuracy is rejected in favor of the IMF forecasts on four occasions (all referring to GDP growth at long horizons), whereas there are no rejections in favor of a benchmark model. These results partly reflect the short time series dimension of our evaluation sample and the associated low power of the test.
@@
Figure 3: Interval scores (IS), separately for (a) inflation and (b) GDP growth, evaluation period 2013-2023. The IS is decomposed into overprediction, underprediction and dispersion, as indicated by the shaded areas. Scores are computed as a weighted sum of interval scores at the $50\%$ and $80\%$ level.
@@
Figure 4: Empirical prediction interval coverage levels, for (a) inflation and (b) GDP growth, evaluation period 2013-2023. In each panel, (I) shows the overall empirical coverage for all methods, with translucent smaller points in the background representing individual coverage for each country-horizon pair, indicating underlying variability in coverage levels. (II) and (III) show coverage levels by forecast horizon, for 50% and 80% nominal coverage, respectively.
@@ 3.5.1 OECD Truth Values
The results in Figures 3 and 4 are based on annual truth values provided by the IMF. By contrast, the benchmark models described in Section 3.3 are fitted to higher-frequency (quarterly) data that are not provided by the IMF, and were thus chosen from another source (OECD). While we made our best effort to choose compatible series, full consistency between the two sources seems unrealistic given the complexity of international macroeconomic aggregates. As a consequence, our use of IMF data for evaluation may pose an unfair disadvantage to the benchmark models that were fitted on OECD data. To examine this possibility, we consider an alternative set of annual truth data that we construct from quarterly OECD data. We use these truth data for both the calculation of forecast errors and the subsequent evaluation. Figure 9 in the supplement studies how this change affects interval scores. The figure focuses on the comparison between the IMF-based forecasts and BVAR-direct as an illustrative benchmark method. In a majority of setups, the different choice of truth value does not change the model ranking between the IMF and BVAR-direct. However, for the shortest horizon (fall, current year), the model ranking is typically reversed in favor of the BVAR-direct method. These results are confirmed by Figures 10 and 11 in the supplement. Thus, while the IMF-based forecasts remain competitive at the other forecast horizons, their good performance at the shortest horizon vanishes when using OECD rather than IMF truth values. While somewhat unsatisfactory, this finding is plausible from a statistical perspective: The shortest forecast horizon is characterized by fairly high predictability since relevant components of the predictand (e.g., GDP growth in 2022) are known when the forecast is issued (e.g., September 2022). In this low-noise setting, the details of the definition of the truth value matter. Hence the IMF-based model outperforms the benchmarks when evaluated against IMF data, whereas the benchmarks prevail when evaluated against OECD data. This effect is not present at longer forecast horizons, where predictability is much lower, so that differences in the definition of the truth value are dominated by other sources of forecast error.
@@ 3.5.2 Additional Intervals and Measures of Forecast Performance
Additionally, Figure 12 in the supplement presents coverage statistics for the corresponding forecast quantiles, where a forecast interval with $\alpha=1-\tau$ induces the forecast quantiles $\hat{q}_{\alpha/2}$ and $\hat{q}_{1-\alpha/2}$. As shown there, the proposed IMF-based method attains close-to-nominal calibration in most instances, whereas some of the benchmark models exhibit considerable deviations from nominal coverage, especially for GDP growth. Nevertheless, for inflation and quantile levels larger than 0.85, some benchmarks, e.g., BVAR-direct, attain better calibration than the IMF-based forecasts.
@@ 3.5.3 Additional Benchmark Models
We find that previous results are in general robust to concrete specifications of the benchmark models. ‘AR-annual’ exhibits high scores on current-year forecasts, where its informational disadvantage is particularly high, but performs comparably to other models on next-year forecasts. In terms of the CRPS, some alternative specifications outperform the respective counterparts from the main analysis on a subset of target-horizon combinations. In particular, the AR($p$) specification outperforms the AR(1) variant for inflation at the next-year horizons. Overall, however, any improvements of alternative specifications over the main benchmarks tend to be small and inconsistent across setups. Compared to our proposed IMF-based approach, the BVAR-Mix model performs slightly better for inflation at both next-year horizons, but performs similar or worse otherwise. This model additionally requires a manual choice of regime switch (necessitated by Covid-19 outliers) which can be viewed as a conceptual drawback as compared to the other models we consider. The ARX-annual method performs worse than the IMF-based approach in most setups. Given that both methods are based on the same information set, this implies that the modeling assumptions of the IMF-based approach are more suitable from a forecasting perspective. Among all benchmarks, the ensemble tends to perform best overall, and similarly well as our proposed IMF-based method.
@@ 3.5.5 Forecasts During the Covid-19 Period
From an ex post perspective, low responsiveness to Covid-related shocks turned out advantageous in the case of GDP growth, where the extreme forecast error in 2020 was followed by moderate forecast errors afterwards. For inflation, where IMF-based intervals were already quite narrow compared to the benchmark models, the low immediate reactiveness of the quantile-based method was a disadvantage, especially if the increase in forecast errors after 2022 represents a continued structural change beyond the end of our study period, which ends in 2023. Within the scope of our prospective study, see Section 3.6, we will be able to monitor the further performance of IMF-based forecast intervals for inflation.
@@ 3.6 Outlook: Prospective Real-Time Forecasts and Evaluation
Initial releases of the forecasts for October 2023 and April 2024, and October 2024 are available in a public GitHub repository (Becker et al., 2023), ensuring transparency and accountability through time stamps. The latter repository also links to the website https://probability-forecasting.shinyapps.io/macropi/ on which we provide concise visualizations of all forecasts and prediction intervals. Moreover, we have deposited a preregistration protocol (Becker et al., 2024) that covers forecasts from October 2024 onwards. In this protocol, we make a commitment regarding how we will evaluate forecasts in 2029, after collecting prospective forecast intervals for five years.
@@ B.2 Individual Interval Score Results
1){1)}1 ) Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the interval score for the $50\%$ and $80\%$ intervals, respectively.
2){2)}2 ) Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.3.2 Evaluation Metrics
GDP Growth Fall, Current 0.23 0.24 0.49 0.43 0.76 0.65
@@
Inflation Fall, Current 0.12 0.12 0.52 0.44 0.76 0.64
@@
1){1)}1 ) A weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$, with weights as given by Bracher et al. (2021a).
2){2)}2 ) Interval coverage at the $50\%$ and $80\%$ levels.
@@ B.4 Choice of Truth Source
[…]

### v4

@@ 3.2 Choice of Tuning Parameters
Table 5 and Figure 8 in the supplement summarize the training sample performance for the different combinations of the tuning parameters. Based on the performance in the training sample, and the considerations in Section 2, we use a rolling window of $R=11$ observations, the absolute error method, and the default empirical quantile method (type = 7) [^1] Note that for $R=11$ and the absolute error method, this is also equivalent to the inverse of the empirical distribution function, that is, type = 1. As we do not estimate quantiles at the tails of the distribution the arguments of Hyndman and Fan (1996) in favor of the type = 8 method do not apply for our setting. of the quantile function in R. Our training sample results indicate that shorter window lengths $R<11$ perform worse in terms of the IS, see Figure 8 in the supplement. The difference in performance is particularly pronounced for very short window lengths $4\leq R<7$ and at longer forecast horizons. By contrast, the benefit of increasing $R$ from $8$ to $11$ is remarkably small in all empirical setups (absolute and directional errors, all forecast horizons). Hence the marginal benefit of increasing the sample size seems to vanish rapidly as $R$ increases. We exclude choices $R>11$ since we cannot evaluate their performance given the time span of our training data, while an extension of the training data would have reduced the time span of our holdout sample, thus limiting our ability to assess out-of-sample performance. From a statistical perspective, using a short window length $R$ ensures adaptability to potential structural breaks (c.f. Inoue et al., 2017). We find that coverage rates on the training data are generally more favorable for the absolute than for the directional error method, sometimes on the order of about 10 percentage points (see Table 5 in the supplement), while interval scores are similar. We additionally present illustrative insights from the training period in Section B.3.1 in the supplement. Albeit attractive in principle, we argue that directional forecast errors sometimes display undesirable practical properties in our setup, and are thus not a good candidate to account for a potential skew in the distribution of forecast errors.
@@ 3.3 Benchmark Models
In order to put the performance of our proposed method in perspective, we compare it to several benchmarks. For ease of presentation, we focus on three main benchmark models in our main empirical analysis that each represent one of the following broader model types: a) univariate or b) multivariate time series models yielding different point predictions, where prediction intervals are calculated either based on past performance forecast errors as in our approach or based on c) the respective time series model specification. For further variants of these benchmark models and respective comparisons we refer to Section 3.5.3 below as well as Sections B.5 and B.6 of the supplement. In particular, Tables 6 and 7 of the supplement provide comprehensive results in terms of evaluation metrics.
All considered benchmarks are of (vector) autoregressive time series form, which is one of the most common choices in macroeconomic forecasting. For example, Faust and Wright (2009) show that simple univariate autoregressions provide good forecasts of U.S. GDP growth when equipped with an appropriate information set, and Carriero et al. (2024a) argue that vector autoregressive models with stochastic volatility compare favorably to nonlinear quantile regression methods for U.S. GDP growth, inflation and unemployment. We fit the main benchmark models to quarterly data from the OECD for inflation and GDP growth (see Section B.1 in the supplement), using expanding windows of training data. In order to roughly match the timing of IMF forecasts, we equip the model-based spring forecasts with data up until the first quarter, and the model-based fall forecasts with data up until the third quarter. As we argue in Section 3.5.1, evaluation with respect to IMF data is still justified.
@@
For the second main benchmark (denoted BVAR), we consider a bivariate Bayesian vector autoregressive specification for inflation and GDP growth. As for the IMF forecasts (and for the AR model), we construct prediction intervals from empirical quantiles of the BVAR’s past forecast errors. Our BVAR implementation is based on the R package bvarsv (Krüger, 2015). While the latter considers the flexible model by Primiceri (2005) and Del Negro and Primiceri (2015), we restrict the priors in our main benchmark to shut off time variation in the model’s mean and covariance parameters, as in Knüppel et al. (2023, Section 5). We find that this simpler specification yields relatively stable parameter estimates for all countries and time periods. By contrast, Primiceri’s original priors that allow for time-varying parameters and stochastic volatility run into problems for some countries such as France and Great Britain after the extreme GDP growth rates observed at the onset of the Covid-19 pandemic. [^2] This problem could possibly be resolved by adjusting the model’s prior parameters to allow for an intermediate degree of flexibility, or by using a stochastic volatility specification that explicitly addresses Covid-19 outliers (as in Carriero et al., 2024b). However, the details of such an adjustment would create many additional degrees of freedom in our setup, given that the severity of the Covid-19 outliers differs across the G7 countries. We hence use a simpler specification with constant parameters for our main analysis. Within the additional analyses discussed in Section 3.5.3, we also consider an approach that uses a stochastic volatility specification until 2019 and a constant volatility specification afterwards. This mixed variant entails a manual choice of regime switch. Even then, however, the performance of the model with constant parameters is largely similar with the exception of inflation at the next-year horizons where the mixed BVAR can outperform all methods and in particular our simple IMF-based approach.
@@ 3.4 Forecast Performance on the Hold-Out Data
We also consider Diebold and Mariano (1995) for the null hypothesis of equal predictive accuracy of the IMF-based method and each benchmark method. In order to limit the number of tests conducted, we average the interval scores across countries at each date (for a given variable, forecast horizon and benchmark model). Pooling information in this way can further provide greater statistical power in empirical applications such as ours (Qu et al., 2024). This yields 24 tests in total (two variables, four horizons, and the AR, BVAR and BVAR-direct benchmark models). We consider two-sided tests, a significance level of 95% and use the R package sandwich and its function NeweyWest (Zeileis, 2004; Zeileis et al., 2020) for the autocorrelation-robust variance estimator that enters the test statistic. This test statistic corresponds to $J_{n,T}^{\mathrm{DM}}$ in Qu et al. (2024) and is a special case of the test statistic $S_{nT}^{3}$ in Akgun et al. (2024). Among the 24 test setups, the null of equal predictive accuracy is rejected in favor of the IMF forecasts on four occasions (all referring to GDP growth at long horizons), whereas there are no rejections in favor of a benchmark model. We further consider $\tilde{S}_{nT}^{3}$ from Akgun et al. (2024), which does not require $T$ to be large but also cannot account for serial correlation in the loss differentials of the test statistic. $\tilde{S}_{nT}^{3}$ relies on the critical values of the Student’s t-distribution which leads to a more conservative result, with only one rejection in favor of the IMF forecasts. [^4] The p-values change from 0.7%, 2.5%, 2.8% and 0.6% to 5.4%, 19.4%, 17.5%, 1.0%, respectively. These results partly reflect the short time series dimension of our evaluation sample and the associated low power of the tests.
@@
Figure 3: Interval scores (IS), separately for (a) inflation and (b) GDP growth, evaluation period 2013-2023. For each forecast horizon, the IS is shown side-by-side for the four methods. The IS is decomposed into underprediction, overprediction and dispersion, as indicated by the vertically stacked shaded areas. Scores are computed as a weighted sum of interval scores at the $50\%$ and $80\%$ level.
@@
Figure 4: Empirical prediction interval coverage levels, for (a) inflation and (b) GDP growth, evaluation period 2013-2023. In each panel, (I) shows the overall empirical coverage for all methods, with translucent smaller points in the background representing individual coverage for each country-horizon pair, indicating underlying variability in coverage levels. (II) and (III) show coverage levels by forecast horizon, for 50% and 80% nominal coverage, respectively. The horizontal lines indicate the respective nominal coverage level.
@@ 3.5.1 OECD Truth Values
The results in Figures 3 and 4 are based on annual truth values provided by the IMF. By contrast, the benchmark models described in Section 3.3 are fitted to higher-frequency (quarterly) data that are not provided by the IMF, and were thus chosen from another source (OECD). While we made our best effort to choose compatible series, full consistency between the two sources seems unrealistic given the complexity of international macroeconomic aggregates. As a consequence, our use of IMF data for evaluation may pose an unfair disadvantage to the benchmark models that were fitted on OECD data. To examine this possibility, we consider an alternative set of annual truth data that we construct from quarterly OECD data. We use these truth data for both the calculation of forecast errors and the subsequent evaluation. Figure 9 in the supplement studies how this change affects interval scores. The figure focuses on the comparison between the IMF-based forecasts and BVAR-direct as an illustrative benchmark method. In a majority of setups, the different choice of truth value does not change the model ranking between the IMF and BVAR-direct. However, for the shortest horizon (“Fall, Current”), the model ranking is typically reversed in favor of the BVAR-direct method. These results are confirmed by Figures 10 and 11 in the supplement. Thus, while the IMF-based forecasts remain competitive at the other forecast horizons, their good performance at the shortest horizon vanishes when using OECD rather than IMF truth values. While somewhat unsatisfactory, this finding is plausible from a statistical perspective: The shortest forecast horizon is characterized by fairly high predictability since relevant components of the predictand (e.g., GDP growth in 2022) are known when the forecast is issued (e.g., September 2022). In this low-noise setting, the details of the definition of the truth value matter. Hence the IMF-based model outperforms the benchmarks when evaluated against IMF data, whereas the benchmarks prevail when evaluated against OECD data. This effect is not present at longer forecast horizons, where predictability is much lower, so that differences in the definition of the truth value are dominated by other sources of forecast error.
@@ 3.5.2 Additional Intervals and Measures of Forecast Performance
Additionally, Figure 12 in the supplement presents coverage statistics for the corresponding forecast quantiles, where a forecast interval with $\alpha=1-\tau$ induces the forecast quantiles $\hat{q}_{\alpha/2}$ and $\hat{q}_{1-\alpha/2}$.
As shown there, the proposed IMF-based method attains close-to-nominal calibration in most instances, whereas some of the benchmark models exhibit considerable deviations from nominal coverage, especially for GDP growth. Nevertheless, for inflation and quantile levels larger than 0.85, some benchmarks, e.g., BVAR-direct, attain better calibration than the IMF-based forecasts.
@@ 3.5.3 Additional Benchmark Models
We find that previous results are in general robust to concrete specifications of the benchmark models. ‘AR-annual’ exhibits high scores on current-year forecasts, where its informational disadvantage is particularly high, but performs comparably to other models on next-year forecasts. In terms of the CRPS, some alternative specifications outperform the respective counterparts from the main analysis on a subset of target-horizon combinations. In particular, the AR($p$) specification outperforms the AR(1) variant for inflation at the next-year horizons. Overall, however, any improvements of alternative specifications over the main benchmarks tend to be small and inconsistent across setups.
Compared to our proposed IMF-based approach, the BVAR-Mix model performs slightly better for inflation at both next-year horizons, but performs similar or worse otherwise. This model additionally requires a manual choice of regime switch (necessitated by Covid-19 outliers) which can be viewed as a conceptual drawback as compared to the other models we consider. The ARX-annual method performs worse than the IMF-based approach in most setups. Given that both methods are based on the same information set, this implies that the modeling assumptions of the IMF-based approach are more suitable from a forecasting perspective.
Among all benchmarks, the ensemble tends to perform best overall, and similarly well as our proposed IMF-based method.
@@ 3.5.5 Forecasts During the Covid-19 Period
From an ex-post perspective, low responsiveness to Covid-19-related shocks turned out advantageous in the case of GDP growth, where the extreme forecast error in 2020 was followed by moderate forecast errors afterwards. For inflation, where IMF-based intervals were already quite narrow compared to the benchmark models, the low immediate reactiveness of the quantile-based method was a disadvantage, especially if the increase in forecast errors after 2022 represents a continued structural change beyond the end of our study period, which ends in 2023. Within the scope of our prospective study, see Section 3.6, we will be able to monitor the further performance of IMF-based forecast intervals for inflation.
@@ 3.6 Outlook: Prospective Real-Time Forecasts and Evaluation
Initial biannual releases of the forecasts starting October 2023 are available in a public GitHub repository (Becker et al., 2023), ensuring transparency and accountability through time stamps. The latter repository also links to the website https://probability-forecasting.shinyapps.io/macropi/ on which we provide concise visualizations of all forecasts and prediction intervals. Moreover, we have deposited a preregistration protocol (Becker et al., 2024) that covers forecasts from October 2024 onwards. In this protocol, we make a commitment regarding how we will evaluate forecasts in 2029, after collecting prospective forecast intervals for five years.
@@ B.2 Individual Interval Score Results
1){1)} Different Versions of the Interval Score. $\text{IS}_{50}$ and $\text{IS}_{80}$ represent the interval score for the $50\%$ and $80\%$ intervals, respectively.
2){2)} Forecast quantiles are taken directly from the parametric forecast distribution of the respective model.
@@ B.3.2 Evaluation Metrics
GDP Growth
Fall, Current 0.23 0.24 0.49 0.43 0.76 0.65
@@
Inflation
Fall, Current 0.12 0.12 0.52 0.44 0.76 0.64
@@
1){1)} A weighted sum of interval scores from quantile forecasts at base levels from main analysis $\tau\in\{0.5,0.8\}$, with weights as given by Bracher et al. (2021a).
2){2)} Interval coverage at the $50\%$ and $80\%$ levels.
@@ B.4 Choice of Truth Source
Figure 9: Difference in interval scores (IS) of IMF-based and BVAR - direct forecasts $\text{IS}(\text{IMF})-\text{IS}(\text{BVAR-direct})$, when using competing truth values as outcome data. Score differences from using IMF truth values are shown on the x axis, analogously for OECD truth values on the y-axis. Each data point corresponds to one combination of horizon, target variable, and country. Values in quadrants “II” and “IV” represent agreement between the two truth sources on which forecasts are better. Values in quadrant “I” represent instances where the IMF truth prefers the IMF forecasts, but the OECD truth prefers the BVAR forecasts. Values in quadrant “IV” represent instances where the IMF truth prefers the BVAR forecasts, but the OECD truth prefers the IMF forecasts.